### v2

@@ 3.2 Conjugate Residual
In this section, we show that under Algorithm 2, CR is able to converge to a normal solution for inconsistent systems (Theorem 3.13). In doing so, we will also demonstrate the similarities between CR and CG, while highlighting additional advantages of CR. Similar to CG, we will then introduce a few additional steps within CR to recover the pseudo-inverse solution (Theorem 3.18). Finally, we will show that, as long as an unlucky breakdown has not occurred, CR generates iterates that are identical to those of MINRES for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$ (Theorem 3.23).
@@ 3.2.1 Convergence Behavior
The convergent behaviors of CR under various conditions have been extensively studied (see Table 2). Originally developed by Hestenes and Stiefel 1, CR was designed for positive definite matrices and can be straightforwardly applied to negative definite matrices. Luenberger 11 later extended CR to indefinite, albeit non-singular, matrices. For inconsistent systems, Hayami and Sugihara 23, 27, 28 established that for inconsistent systems, the necessary and sufficient condition for CR to avoid unlucky breakdown is that the matrix $\mathbf{A}$ is positive semi-definite. In particular, they demonstrated that for any indefinite $\mathbf{A}$, there exists a right-hand side vector $\mathbf{b}$ that will trigger an unlucky breakdown. While this is a crucial insight, it remains of interest to investigate the behavior of CR for the many other vectors $\mathbf{b}$ that do not cause this breakdown. We do that here. Specifically, we show that for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$, if no unlucky breakdown occurs, CR will indeed converge to a normal solution. Furthermore, we devise a few extra steps for CR to obtain the pseudo-inverse solution.
@@
(Semi-)Definite Hestenes & Stiefel 1 Hayami & Sugihara 23, 27, 28
Indefinite Luenberger (non-singular matrix) 11 Theorems 3.13 and 3.18
Table 2: Relevant studies on the convergent behaviors of CR under various conditions.
Remark 3.10.
There are many algorithms considering the alternative Krylov subspaces, like $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ or $\mathcal{K}(\mathbf{A^{\top}A},\mathbf{Ab})$, instead of $\mathcal{K}(\mathbf{A},\mathbf{b})$, for example MR-II / RRMR 29 and variants of CG and CR 30, 31. The advantages of these alternative Krylov subspaces are that the vector $\mathbf{Ab}$ is in the range of $\mathbf{A}$ and, hence, avoids unnecessary complications arising from the inconsistent systems. However, it is also empirically shown that these alternative subspaces may not produce better results than $\mathcal{K}(\mathbf{A},\mathbf{b})$ and has more computational costs, see the work 32 discussing MINRES and RRMR. It is also shown that the iterates of Krylov subspace methods defined in $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ tend to have a slower convergence than $\mathcal{K}(\mathbf{A},\mathbf{b})$, see the work33 comparing SYMMLQ and CG.
One of the similarities between CR and CG is that, as in Lemma 3.3 for CG, at termination for CR we also have $\mathbf{p}_{g-1}\in\text{Null}(\mathbf{A})$.
Lemma 3.11.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Algorithm 2, in Algorithm 2, we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof 3.12.
The proof follows a similar line of reasoning as that of Lemma 3.3. Let $k=g$. If $g=1$, since $\mathbf{b}\notin\text{Range}(\mathbf{A})$, we must have $\mathbf{p}_{0}=\mathbf{b}\in\text{Null}(\mathbf{A})$ and the claim holds. So consider $g>1$. By Section 3-(a) and Definition 1.2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
@@
In addition, from Section 3-(c), we have $\left\langle\mathbf{A}\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle$. We note that $\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle\neq 0$ by Algorithm 2. Hence, we must have $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
@@
where the last equality follows from Section 3-(c). Again, since $\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{r}_{i-1}\right\rangle\neq 0$, by Algorithm 2, it follows that $c_{i-1}=0$. As a result, we must have got $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, the constant $\alpha_{g-1}=\left\langle\mathbf{r}_{g-1},\mathbf{A}\mathbf{r}_{g-1}\right\rangle/\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$ can be, in principle, calculated in the next iterate. We first note that by Section 3-(c), we have
@@
Conversely, let $1\leq k\leq g$ be the first iteration where $\mathbf{Ap}_{k-1}=\mathbf{0}$. We first show that $\mathbf{p}_{k-1}\neq\mathbf{0}$. If $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. Suppose $k\geq 2$ and $\mathbf{p}_{k-1}=\mathbf{0}$. Clearly, $\beta_{k-2}\neq 0$ as otherwise $\mathbf{0}=\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ also implies $\mathbf{r}_{k-1}=\mathbf{0}$ which by assumption on $\mathbf{b}$ is impossible. However, by Section 3-(a), $\mathbf{r}_{k-1}$ and $\mathbf{p}_{k-2}$ must be linearly independent and we arrive at a contradiction. As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$.
@@
While Lemmas 3.3 and 3.11 highlight a similarity between CG and CR, Theorems 3.1 and 3.13 showcase a drastic difference. Specifically, we now show that, in sharp contrast to CG, CR always recovers a normal solution.
Theorem 3.13 (CR’s Ability to Recover Normal Solution).
@@
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Algorithm 2, in Algorithm 2 we have $\mathbf{Ar}_{k-1}=\mathbf{0}$ if and only if $k=g$.
Proof 3.14.
@@
Suppose $k=g$. From $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}$ and Lemma 3.11, we have $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle+\beta_{g-2}\left\langle\mathbf{r}_{g-1},\mathbf{Ap}_{g-2}\right\rangle=0$, which using Section 3-(c) gives $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle=0$, implying $\beta_{g-2}=0$ by its definition in Algorithm 2. Hence, $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$ and the result follows from Lemma 3.11.
@@
Conversely, suppose $\mathbf{Ar}_{k-1}=\mathbf{0}$ for some iteration $1\leq k\leq g$. From $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$, we get $\mathbf{Ap}_{k-1}=\beta_{k-2}\mathbf{Ap}_{k-2}$, which using Section 3-(c) gives $\|\mathbf{Ap}_{k-1}\|^{2}=\beta_{k-2}\left\langle\mathbf{Ap}_{k-1},\mathbf{Ap}_{k-2}\right\rangle=0$. Hence, $\mathbf{Ap}_{k-1}=\mathbf{0}$, which by Lemma 3.11 can only happen when $k=g$.
Remark 3.15.
@@
For both Algorithms 1 and 2, we always have $\mathbf{p}_{0}=\mathbf{r}_{0}$. However, from the proof of Theorem 3.13, it can be seen that under the assumptions of Theorem 3.13, in Algorithm 2, we also have $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$. It is important to note that this property is specific to CR, and CG does not exhibit the same characteristics.
@@
While, by Theorem 3.13, CR is guaranteed to converge to a normal solution, we next show that we can in fact obtain the pseudo-inverse solution from the last iterate of CR through a straightforward procedure. This stands in sharp contrast to CG, where a similar procedure requires non-trivial modifications to the original algorithm. To achieve this, we first establish several technical lemmas.
Lemma 3.16.
@@
We note that, as alluded to in Remark 3.5, both procedures CR$(\mathbf{A},\tilde{\mathbf{b}})$ and CR$(\mathbf{A},\mathbf{b})$ undergo precisely the same number of iterations. Hence, the assertion of Lemma 3.16 holds true for all iterations in both procedures.
Since $\tilde{\mathbf{b}}\in\text{Range}(\mathbf{A})$, by Fact 1, for the procedure CR($\mathbf{A},\tilde{\mathbf{b}}$), we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{g-1}=\tilde{\mathbf{x}}_{g-1}=\tilde{\mathbf{x}}_{\tilde{g}}=\mathbf{A}^{\dagger}\tilde{\mathbf{b}}=\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{x}^{+}$. Now, using a similar line of reasoning as that leading up to Theorem 3.8 for CG, by utilizing Lemma B.1, for some constant $c$, we have
@@
which using Lemma 3.11 gives $c=\left\langle\mathbf{p}_{g-1},\mathbf{x}_{g-1}\right\rangle/\|\mathbf{p}_{g-1}\|^{2}$. Hence,
@@
which leads us to Theorems 3.18 and 4.
@@
Theorem 3.18 (Pseudo-inverse Solution with Algorithm 4).
Under Algorithm 2, in Algorithm 4, we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
@@
Under Algorithm 2, we simply have
@@
which implies that the norm of the residuals in Algorithm 2 (or Algorithm 4) monotonically decreasing. On the other hand, MINRES also guarantees a monotonic decrease of the residual norms. The equivalence between CR and MINRES for PD systems is well-studied 21. Next, we show that, in the absence of an unlucky breakdown, this equivalence indeed extends beyond PD settings to all systems.

### v4

@@ 3.2 Conjugate Residual
In this section, we show that under Assumption 1, CR is able to converge to a normal solution for inconsistent systems (Theorem 3). In doing so, we will also demonstrate the similarities between CR and CG, while highlighting additional advantages of CR. Similar to CG, we will then introduce a few additional steps within CR to recover the pseudo-inverse solution (Theorem 4). Finally, we will show that, as long as an unlucky breakdown has not occurred, CR generates iterates that are identical to those of MINRES for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$ (Theorem 5).
@@ 3.2.1 Convergence Behavior
The convergent behavior of CR under various conditions have been extensively studied (see Table 2). Originally developed by Hestenes and Stiefel [22], CR was designed for positive definite matrices and can be straightforwardly applied to negative definite matrices. Luenberger [30] later extended CR to indefinite, albeit non-singular, matrices. For inconsistent systems, Hayami and Sugihara [18, 21, 20] established that for inconsistent systems, the necessary and sufficient condition for CR to avoid unlucky breakdown is that the matrix $\mathbf{A}$ is positive semi-definite. In particular, they demonstrated that for any indefinite $\mathbf{A}$, there exists a right-hand side vector $\mathbf{b}$ that will trigger an unlucky breakdown. While this is a crucial insight, it remains of interest to investigate the behavior of CR for the many other vectors $\mathbf{b}$ that do not cause this breakdown. We do that here. Specifically, we show that for any symmetric matrix $\mathbf{A}$ and any vector $\mathbf{b}$, if no unlucky breakdown occurs, CR will indeed converge to a normal solution. Furthermore, we devise a few extra steps for CR to obtain the pseudo-inverse solution.
@@
(Semi-)Definite Hestenes & Stiefel [22] Hayami & Sugihara [18, 21, 20]
Indefinite Luenberger [30] & Greenbaum [15] Theorems 3 and 4
Table 2: Relevant studies on the convergent behavior of CR under various conditions.
Remark 7.
There are many algorithms considering the alternative Krylov subspaces, like $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ or $\mathcal{K}(\mathbf{A^{\top}A},\mathbf{Ab})$, instead of $\mathcal{K}(\mathbf{A},\mathbf{b})$, for example MR-II / RRMR [16] and variants of CG and CR [2, 11]. The advantages of these alternative Krylov subspaces are that the vector $\mathbf{Ab}$ is in the range of $\mathbf{A}$ and, hence, avoids unnecessary complications arising from the inconsistent systems. However, it is also empirically shown that these alternative subspaces may not produce better results than $\mathcal{K}(\mathbf{A},\mathbf{b})$ and has more computational costs, see the work [3] discussing MINRES and RRMR. It is also shown that the iterates of Krylov subspace methods defined in $\mathcal{K}(\mathbf{A},\mathbf{Ab})$ tend to have a slower convergence than $\mathcal{K}(\mathbf{A},\mathbf{b})$, see the work[7] comparing SYMMLQ and CG.
One of the similarities between CR and CG is that, as in Lemma 1 for CG, at termination for CR we also have $\mathbf{p}_{g-1}\in\text{Null}(\mathbf{A})$.
Lemma 3.
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 2, we have $\mathbf{Ap}_{k-1}=\mathbf{0}$ if and only if $k=g$. Furthermore, $\mathbf{p}_{g-1}\neq\mathbf{0}$.
Proof.
The proof follows a similar line of reasoning as that of Lemma 1. Let $k=g$. If $g=1$, since $\mathbf{b}\notin\text{Range}(\mathbf{A})$, we must have $\mathbf{p}_{0}=\mathbf{b}\in\text{Null}(\mathbf{A})$ and the claim holds. So consider $g>1$. By Property 1-(a) and Definition 2, we note that $\mathbf{Ap}_{g-1}\in\mathcal{K}_{g}(\mathbf{A},\mathbf{b})$ and
@@
In addition, from Property 1-(c), we have $\left\langle\mathbf{A}\mathbf{p}_{i-1},\mathbf{Ap}_{g-1}\right\rangle=0$, for all $11\leq i<g-$ 1. For $i=1$, $0=\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{Ap}_{g-1}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{p}_{0},\mathbf{r}_{0}\right\rangle=c_{0}\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle$. We note that $\left\langle\mathbf{A}\mathbf{r}_{0},\mathbf{r}_{0}\right\rangle\neq 0$ by Assumption 1. Hence, we must have $c_{0}=0$. Similarly, for $12\leq i<g-$ 1,
@@
where the last equality follows from Property 1-(c). Again, since $\left\langle\mathbf{A}\mathbf{r}_{i-1},\mathbf{r}_{i-1}\right\rangle\neq 0$, by Assumption 1, it follows that $c_{i-1}=0$. As a result, we must have got $\mathbf{Ap}_{g-1}=c_{g-1}\mathbf{r}_{g-1}$. Now, if $c_{g-1}=0$, then we are done. Otherwise, assuming $\mathbf{Ap}_{g-1}\not=\mathbf{0}$, since $\mathbf{r}_{g-1}\neq\mathbf{0}$, the constant $\alpha_{g-1}=\left\langle\mathbf{r}_{g-1},\mathbf{A}\mathbf{r}_{g-1}\right\rangle/\left\langle\mathbf{A}\mathbf{p}_{g-1},\mathbf{Ap}_{g-1}\right\rangle$ can be, in principle, calculated in the next iterate. We first note that by Property 1-(c), we have
@@
Conversely, let $1\leq k\leq g$ be the first iteration where $\mathbf{Ap}_{k-1}=\mathbf{0}$. We first show that $\mathbf{p}_{k-1}\neq\mathbf{0}$. If $k=1$, we have $\mathbf{p}_{0}=\mathbf{b}\neq\mathbf{0}$. Suppose $k\geq 2$ and $\mathbf{p}_{k-1}=\mathbf{0}$. Clearly, $\beta_{k-2}\neq 0$ as otherwise $\mathbf{0}=\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$ also implies $\mathbf{r}_{k-1}=\mathbf{0}$ which by assumption on $\mathbf{b}$ is impossible. However, by Property 1-(a), $\mathbf{r}_{k-1}$ and $\mathbf{p}_{k-2}$ must be linearly independent and we arrive at a contradiction. As a result, since $\mathbf{p}_{k-1}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ and $\mathbf{p}_{k-1}\neq\mathbf{0}$, we can write $\mathbf{p}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i}\mathbf{b}$ and at least one $c_{j}$ must be non-zero. Hence, $\mathbf{0}=\mathbf{Ap}_{k-1}=\sum_{i=0}^{k-1}c_{i}\mathbf{A}^{i+1}\mathbf{b}$, which implies $\dim(\mathcal{K}_{k+1}(\mathbf{A},\mathbf{b}))=\dim(\mathcal{K}_{k}(\mathbf{A},\mathbf{b}))$, and so we must have $k=g$. ∎
@@
While Lemmas 1 and 3 highlight a similarity between CG and CR, Theorems 1 and 3 showcase a drastic difference. Specifically, we now show that, in sharp contrast to CG, CR always recovers a normal solution.
Theorem 3 (CR’s Ability to Recover Normal Solution).
@@
Let $\mathbf{b}\notin\text{Range}(\mathbf{A})$. Under Assumption 1, in Algorithm 2 we have $\mathbf{Ar}_{k-1}=\mathbf{0}$ if and only if $k=g$.
Proof.
@@
Suppose $k=g$. From $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}+\beta_{g-2}\mathbf{p}_{g-2}$ and Lemma 3, we have $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle+\beta_{g-2}\left\langle\mathbf{r}_{g-1},\mathbf{Ap}_{g-2}\right\rangle=0$, which using Property 1-(c) gives $\left\langle\mathbf{r}_{g-1},\mathbf{Ar}_{g-1}\right\rangle=0$, implying $\beta_{g-2}=0$ by its definition in Algorithm 2. Hence, $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$ and the result follows from Lemma 3.
@@
Conversely, suppose $\mathbf{Ar}_{k-1}=\mathbf{0}$ for some iteration $1\leq k\leq g$. From $\mathbf{p}_{k-1}=\mathbf{r}_{k-1}+\beta_{k-2}\mathbf{p}_{k-2}$, we get $\mathbf{Ap}_{k-1}=\beta_{k-2}\mathbf{Ap}_{k-2}$, which using Property 1-(c) gives $\|\mathbf{Ap}_{k-1}\|^{2}=\beta_{k-2}\left\langle\mathbf{Ap}_{k-1},\mathbf{Ap}_{k-2}\right\rangle=0$. Hence, $\mathbf{Ap}_{k-1}=\mathbf{0}$, which by Lemma 3 can only happen when $k=g$. ∎
Remark 8.
@@
For both Algorithms 1 and 2, we always have $\mathbf{p}_{0}=\mathbf{r}_{0}$. However, from the proof of Theorem 3, it can be seen that under the assumptions of Theorem 3, in Algorithm 2, we also have $\mathbf{p}_{g-1}=\mathbf{r}_{g-1}$. It is important to note that this property is specific to CR, and CG does not exhibit the same characteristics.
@@
While, by Theorem 3, CR is guaranteed to converge to a normal solution, we next show that we can in fact obtain the pseudo-inverse solution from the last iterate of CR through a straightforward procedure. This stands in sharp contrast to CG, where a similar procedure requires non-trivial modifications to the original algorithm. To achieve this, we first establish several technical lemmas.
Lemma 4.
@@
We note that, as alluded to in Remark 5, both procedures CR$(\mathbf{A},\tilde{\mathbf{b}})$ and CR$(\mathbf{A},\mathbf{b})$ undergo precisely the same number of iterations. Hence, the assertion of Lemma 4 holds true for all iterations in both procedures.
Since $\tilde{\mathbf{b}}\in\text{Range}(\mathbf{A})$, by Fact 1, for the procedure CR($\mathbf{A},\tilde{\mathbf{b}}$), we have $\mathbf{A}^{\dagger}\mathbf{A}\mathbf{x}_{g-1}=\tilde{\mathbf{x}}_{g-1}=\tilde{\mathbf{x}}_{\tilde{g}}=\mathbf{A}^{\dagger}\tilde{\mathbf{b}}=\mathbf{A}^{\dagger}\mathbf{b}=\mathbf{x}^{+}$. Now, using a similar line of reasoning as that leading up to Theorem 2 for CG, by utilizing Lemma 12, for some constant $c$, we have
@@
which using Lemma 3 gives $c=\left\langle\mathbf{p}_{g-1},\mathbf{x}_{g-1}\right\rangle/\|\mathbf{p}_{g-1}\|^{2}$. Hence,
@@
which leads us to Theorems 4 and 4.
@@
Theorem 4 (Pseudo-inverse Solution with Algorithm 4).
Under Assumption 1, in Algorithm 4, we have $\mathbf{x}^{+}=\mathbf{A}^{\dagger}\mathbf{b}$.
@@
Under Assumption 1, we simply have
@@
which implies that the norm of the residuals in Algorithm 2 (or Algorithm 4) monotonically decreasing. On the other hand, MINRES also guarantees a monotonic decrease of the residual norms. The equivalence between CR and MINRES for PD systems is well-studied [14]. Next, we show that, in the absence of an unlucky breakdown, this equivalence indeed extends beyond PD settings to all systems.